# An AI Architecture with the Capability to Explain Recognition Results

### v1

@@ II Related Work
Figure 2: A letter ‘S’ example.
TABLE III: Explainability Matrix based on the Recall Metric
PDF Effectiveness PDF Explainability
$F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$
$F_{1}$ Stroke S 0.9229 1.0
$F_{2}$ Circle 1 0.9954 1.0
$F_{3}$ Crossing 1 0.9850 1.0
$F_{4}$ Ellipse n 0.1079 1.0
$F_{5}$ Ell-Cir 1 0.9954 1.0
$F_{6}$ Endpoint S 0.4963 1.0
$F_{7}$ Encl. Reg. 1 0.9996 1.0
$F_{8}$ Line S 0.4950 1.0
$F_{9}$ Convex Hull E 0.2604 1.0
$F_{10}$ Corner S 0.5088 1.0
$F_{11}$ No Property S 0.9238 0.0
Weights $3.975$ $0.2604$ $3.346$ $0.1079$ $3.975$ $0.2604$ $2.423$ $0.1079$
Confidence / Explainability $51.7\%$ $3.39\%$ $43.5\%$ $1.40\%$ $100\%$ $100\%$ $72.4\%$ $100\%$
TABLE IV: Explanation based on the Recall Metric
Class Confidence Explainability Explainable Description
1 51.7% 100% Confidence is medium for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties.
S 43.5% 72.4% Confidence is medium for interpreting this character as an S due to the stroke, corner, endpoint, and line properties.
E 3.39% 100% Confidence is low for interpreting this character as an E due to the convex hull property.
n 1.40% 100% Confidence is low for interpreting this character as an n due to the ellipse property.
@@ III-C Effectiveness Metrics
TABLE V: Explainability Matrix based on the Accuracy Metric
PDF Effectiveness PDF Explainability
$F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$
$F_{1}$ Stroke S 0.9964 1.0
$F_{2}$ Circle 1 0.8417 1.0
$F_{3}$ Crossing 1 0.7354 1.0
$F_{4}$ Ellipse n 0.9774 1.0
$F_{5}$ Ell-Cir 1 0.8444 1.0
$F_{6}$ Endpoint S 0.9776 1.0
$F_{7}$ Encl. Reg. 1 0.3615 1.0
$F_{8}$ Line S 0.9787 1.0
$F_{9}$ Convex Hull E 0.9767 1.0
$F_{10}$ Corner S 0.9791 1.0
$F_{11}$ No Property S 0.9964 0.0
Weights $2.783$ $0.9767$ $4.928$ $0.9774$ $2.783$ $0.9767$ $3.932$ $0.9774$
Confidence / Explainability $28.8\%$ $10.1\%$ $51.0\%$ $10.1\%$ $100\%$ $100\%$ $79.8\%$ $100\%$
TABLE VI: Explanations based on the Accuracy Metric
Class Confidence Explainability Explainable Description
S 51.0% 79.8% Confidence is medium for interpreting this character as an S due to the stroke, corner, endpoint, and line properties.
1 28.8% 100% Confidence is medium for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties.
n 10.1% 100% Confidence is low for interpreting this character as an n due to the ellipse property.
E 10.1% 100% Confidence is low for interpreting this character as an E due to the convex hull property.
@@ IV-B Effectiveness Results
TABLE VII: Explainability Matrix based on the $E_{PARS}$ Metric
PDF Effectiveness PDF Explainability
$F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$
$F_{1}$ Stroke S 0.8341 1.0
$F_{2}$ Circle 1 0.0830 1.0
$F_{3}$ Crossing 1 0.0389 1.0
$F_{4}$ Ellipse n 0.0406 1.0
$F_{5}$ Ell-Cir 1 0.0848 1.0
$F_{6}$ Endpoint S 0.2274 1.0
$F_{7}$ Encl. Reg. 1 0.0040 1.0
$F_{8}$ Line S 0.2401 1.0
$F_{9}$ Convex Hull E 0.1067 1.0
$F_{10}$ Corner S 0.2508 1.0
$F_{11}$ No Property S 0.8385 0.0
Weights $0.2106$ $0.1067$ $2.391$ $0.0406$ $0.2106$ $0.1067$ $1.5524$ $0.0406$
Confidence / Explainability $7.66\%$ $3.88\%$ 86.9% $1.48\%$ $100\%$ $100\%$ 64.9% $100\%$
TABLE VIII: Explanations based on the $E_{PARS}$ Metric
Class Confidence Explainability Explainable Description
S 86.9% 64.9% Confidence is high for interpreting this character as an S due to the stroke, corner, endpoint, and line properties.
1 7.66% 100% Confidence is low for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties.
E 3.88% 100% Confidence is low for interpreting this character as an E due to the convex hull property.
n 1.40% 100% Confidence is low for interpreting this character as an n due to the ellipse property.
@@ IV-C Explainable Example
Fig. 2 depicts the example, an EMNIST balanced test, handwritten letter ‘S’ sample (index 18) labeled as a capital ‘S’. The explainable architecture used for the example is composed of eleven PDFs with SVM IEs. Data from multiple effectiveness metrics (such as Recall, Accuracy, and $E_{PARS}$) will be presented to demonstrate how the explainable system benefits from more robust effectiveness metrics.
The results of processing the example using the Recall metric are in Tables III and IV, the results of using the Accuracy metric are depicted in Tables V and VI, and the results of using the $E_{PARS}$ metric are in Tables VII and VIII. Table IX illustrates the metric particulars for the PDFs in the example in one place. Note that four classes were voted on for the example. Each PDF’s vote and explainability metric is the same across the various tables. Only the effectiveness varies.
When examining the Explainability data from Tables III, V, and VII, each flow identifier, $F_{j}$, is in the first column where $j$ denotes the flow number. The second column indicates the property name. Flow $F_{11}$ is labeled as no property because it represents an unexplainable flow, without an explainable property. The Class Vote column indicates the class that was selected by each flow.
@@
Ordered explanation results are represented in Tables IV, VI, and VIII. The first column indicates the class, $d$. The second column the Confidence. The third column, represents the explainability associated with the decision. The final column is the rationale provided to the user.
Observe in Tables III and IV that the four PDFs that voted for class one had high Recall, reflected in the Effectiveness column $E_{j,1}$ of Table III in bold. This resulted in the digit one winning with medium confidence, $51.7\%$, due to the enclosed region, circle, ellipse-circle, and crossing properties. The second choice was the ‘S’ with medium confidence, $43.5\%$. Since Recall in (3) has $TP$ in the numerator, high $TP$ and very low $FN$ counts of the PDFs that voted for the digit one observed in Table IX explains why the digit one wins using the Recall metric.
Tables V and VI contain data and results from using the Accuracy metric on the example. Accuracy is given in (1). Due to the comparatively high $TN$ rate of the PDFs that voted for the ‘S’, as shown in bold in Table IX it wins with medium confidence, $51.0\%$, due to the stroke, corner, endpoint, and line properties. The ‘S’ also received a vote from the unexplainable PDF. The lower explainability metric, at $79.8\%$, reflects diminished explainability due to no explainable property associated with part ($20.2\%$) of the contribution for the decision.
TABLE IX: Ex. 2 Effectiveness Metrics (%) based on Votes
Property Class TP TN FP FN R ACC $E_{PARS}$
Encl. Reg. 1 2.13 34.0 63.8 0.0 100 36.2 0.40
Circle 1 2.12 82.1 15.8 0.0 99.5 84.2 8.29
Ell-Cir 1 2.12 82.3 15.5 0.0 99.5 84.5 8.47
Crossing 1 2.01 71.5 26.4 0.0 98.5 73.5 3.88
Convex Hull E 0.55 97.1 0.76 1.57 26.0 97.7 10.7
No Property S 1.97 97.7 0.19 0.16 92.4 99.7 83.9
Stroke S 1.96 97.7 0.20 0.16 92.3 99.6 83.4
Corner S 1.08 96.8 1.04 1.05 50.9 97.9 25.1
Endpoint S 1.06 96.7 0.20 0.20 49.6 97.8 22.7
Line S 1.05 96.8 1.05 1.07 49.5 97.9 24.0
Ellipse n 0.23 97.5 0.36 1.90 10.8 97.7 4.07
The final set of Explainability and Explanations Tables, VII and VIII, involve the $E_{PARS}$ metric. The stroke and unexplainable flows have a much higher effectiveness than other metrics. High effectiveness from the Accuracy metric for the flows voting for the digit one were due to insensitivity of Accuracy to $FP$s. The same flows voting for the digit one have a reduced effectiveness due to relative sensitivity of $E_{PARS}$ to $FP$ counts, as observed in Table IX. The capital ‘S’  correctly wins with high confidence, $86.9\%$, using $E_{PARS}$.

### v2

@@ II Related Work
[…]
figureA letter ‘S’ example.
[…]
tableExplainability Matrix based on the Recall Metric PDF Effectiveness PDF Explainability $F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$ $F_{1}$ Stroke S 0.9229 1.0 $F_{2}$ Circle 1 0.9954 1.0 $F_{3}$ Crossing 1 0.9850 1.0 $F_{4}$ Ellipse n 0.1079 1.0 $F_{5}$ Ell-Cir 1 0.9954 1.0 $F_{6}$ Endpoint S 0.4963 1.0 $F_{7}$ Encl. Reg. 1 0.9996 1.0 $F_{8}$ Line S 0.4950 1.0 $F_{9}$ Convex Hull E 0.2604 1.0 $F_{10}$ Corner S 0.5088 1.0 $F_{11}$ No Property S 0.9238 0.0 Weights $3.975$ $0.2604$ $3.346$ $0.1079$ $3.975$ $0.2604$ $2.423$ $0.1079$ Confidence / Explainability $51.7\%$ $3.39\%$ $43.5\%$ $1.40\%$ $100\%$ $100\%$ $72.4\%$ $100\%$
[…]
tableExplanation based on the Recall Metric Class Confidence Explainability Explainable Description 1 51.7% 100% Confidence is medium for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties. S 43.5% 72.4% Confidence is medium for interpreting this character as an S due to the stroke, corner, endpoint, and line properties. E 3.39% 100% Confidence is low for interpreting this character as an E due to the convex hull property. n 1.40% 100% Confidence is low for interpreting this character as an n due to the ellipse property.
@@ III-C Effectiveness Metrics
[…]
tableExplainability Matrix based on the Accuracy Metric PDF Effectiveness PDF Explainability $F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$ $F_{1}$ Stroke S 0.9964 1.0 $F_{2}$ Circle 1 0.8417 1.0 $F_{3}$ Crossing 1 0.7354 1.0 $F_{4}$ Ellipse n 0.9774 1.0 $F_{5}$ Ell-Cir 1 0.8444 1.0 $F_{6}$ Endpoint S 0.9776 1.0 $F_{7}$ Encl. Reg. 1 0.3615 1.0 $F_{8}$ Line S 0.9787 1.0 $F_{9}$ Convex Hull E 0.9767 1.0 $F_{10}$ Corner S 0.9791 1.0 $F_{11}$ No Property S 0.9964 0.0 Weights $2.783$ $0.9767$ $4.928$ $0.9774$ $2.783$ $0.9767$ $3.932$ $0.9774$ Confidence / Explainability $28.8\%$ $10.1\%$ $51.0\%$ $10.1\%$ $100\%$ $100\%$ $79.8\%$ $100\%$
[…]
tableExplanations based on the Accuracy Metric Class Confidence Explainability Explainable Description S 51.0% 79.8% Confidence is medium for interpreting this character as an S due to the stroke, corner, endpoint, and line properties. 1 28.8% 100% Confidence is medium for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties. n 10.1% 100% Confidence is low for interpreting this character as an n due to the ellipse property. E 10.1% 100% Confidence is low for interpreting this character as an E due to the convex hull property.
@@ IV-B Effectiveness Results
[…]
tableExplainability Matrix based on the $E_{PARS}$ Metric PDF Effectiveness PDF Explainability $F_{j}$ Property Class Vote $E_{j,1}$ $E_{j,E}$ $E_{j,S}$ $E_{j,n}$ $Ex_{1}$ $Ex_{E}$ $Ex_{S}$ $Ex_{n}$ $F_{1}$ Stroke S 0.8341 1.0 $F_{2}$ Circle 1 0.0830 1.0 $F_{3}$ Crossing 1 0.0389 1.0 $F_{4}$ Ellipse n 0.0406 1.0 $F_{5}$ Ell-Cir 1 0.0848 1.0 $F_{6}$ Endpoint S 0.2274 1.0 $F_{7}$ Encl. Reg. 1 0.0040 1.0 $F_{8}$ Line S 0.2401 1.0 $F_{9}$ Convex Hull E 0.1067 1.0 $F_{10}$ Corner S 0.2508 1.0 $F_{11}$ No Property S 0.8385 0.0 Weights $0.2106$ $0.1067$ $2.391$ $0.0406$ $0.2106$ $0.1067$ $1.5524$ $0.0406$ Confidence / Explainability $7.66\%$ $3.88\%$ 86.9% $1.48\%$ $100\%$ $100\%$ 64.9% $100\%$
[…]
tableExplanations based on the $E_{PARS}$ Metric Class Confidence Explainability Explainable Description S 86.9% 64.9% Confidence is high for interpreting this character as an S due to the stroke, corner, endpoint, and line properties. 1 7.66% 100% Confidence is low for interpreting this character as a one due to the enclosed region, circle, ellipse-circle, and crossing properties. E 3.88% 100% Confidence is low for interpreting this character as an E due to the convex hull property. n 1.40% 100% Confidence is low for interpreting this character as an n due to the ellipse property.
@@ IV-C Explainable Example
Fig. II depicts the example, an EMNIST balanced test, handwritten letter ‘S’ sample (index 18) labeled as a capital ‘S’. The explainable architecture used for the example is composed of eleven PDFs with SVM IEs. Data from multiple effectiveness metrics (such as Recall, Accuracy, and $E_{PARS}$) will be presented to demonstrate how the explainable system benefits from more robust effectiveness metrics.
The results of processing the example using the Recall metric are in Tables II and II, the results of using the Accuracy metric are depicted in Tables III-C and III-C, and the results of using the $E_{PARS}$ metric are in Tables IV-B and IV-B. Table IV-C illustrates the metric particulars for the PDFs in the example in one place. Note that four classes were voted on for the example. Each PDF’s vote and explainability metric is the same across the various tables. Only the effectiveness varies.
When examining the Explainability data from Tables II, III-C, and IV-B, each flow identifier, $F_{j}$, is in the first column where $j$ denotes the flow number. The second column indicates the property name. Flow $F_{11}$ is labeled as no property because it represents an unexplainable flow, without an explainable property. The Class Vote column indicates the class that was selected by each flow.
@@
Ordered explanation results are represented in Tables II, III-C, and IV-B. The first column indicates the class, $d$. The second column the Confidence. The third column, represents the explainability associated with the decision. The final column is the rationale provided to the user.
Observe in Tables II and II that the four PDFs that voted for class one had high Recall, reflected in the Effectiveness column $E_{j,1}$ of Table II in bold. This resulted in the digit one winning with medium confidence, $51.7\%$, due to the enclosed region, circle, ellipse-circle, and crossing properties. The second choice was the ‘S’ with medium confidence, $43.5\%$. Since Recall in (3) has $TP$ in the numerator, high $TP$ and very low $FN$ counts of the PDFs that voted for the digit one observed in Table IV-C explains why the digit one wins using the Recall metric.
Tables III-C and III-C contain data and results from using the Accuracy metric on the example. Accuracy is given in (1). Due to the comparatively high $TN$ rate of the PDFs that voted for the ‘S’, as shown in bold in Table IV-C it wins with medium confidence, $51.0\%$, due to the stroke, corner, endpoint, and line properties. The ‘S’ also received a vote from the unexplainable PDF. The lower explainability metric, at $79.8\%$, reflects diminished explainability due to no explainable property associated with part ($20.2\%$) of the contribution for the decision.
[…]
tableEx. 2 Effectiveness Metrics (%) based on Votes Property Class TP TN FP FN R ACC $E_{PARS}$ Encl. Reg. 1 2.13 34.0 63.8 0.0 100 36.2 0.40 Circle 1 2.12 82.1 15.8 0.0 99.5 84.2 8.29 Ell-Cir 1 2.12 82.3 15.5 0.0 99.5 84.5 8.47 Crossing 1 2.01 71.5 26.4 0.0 98.5 73.5 3.88 Convex Hull E 0.55 97.1 0.76 1.57 26.0 97.7 10.7 No Property S 1.97 97.7 0.19 0.16 92.4 99.7 83.9 Stroke S 1.96 97.7 0.20 0.16 92.3 99.6 83.4 Corner S 1.08 96.8 1.04 1.05 50.9 97.9 25.1 Endpoint S 1.06 96.7 0.20 0.20 49.6 97.8 22.7 Line S 1.05 96.8 1.05 1.07 49.5 97.9 24.0 Ellipse n 0.23 97.5 0.36 1.90 10.8 97.7 4.07
The final set of Explainability and Explanations Tables, IV-B and IV-B, involve the $E_{PARS}$ metric. The stroke and unexplainable flows have a much higher effectiveness than other metrics. High effectiveness from the Accuracy metric for the flows voting for the digit one were due to insensitivity of Accuracy to $FP$s. The same flows voting for the digit one have a reduced effectiveness due to relative sensitivity of $E_{PARS}$ to $FP$ counts, as observed in Table IV-C. The capital ‘S’  correctly wins with high confidence, $86.9\%$, using $E_{PARS}$.